To pay for your education, youve taken out $18,000 in student loans. H you make monthiy payments over 12 years at 8 percent compounded monthly, how much are your monthly student loan payments? Your monthly student loan payments are (Round to the nearest cent.)
Answer #1

PV = 18,000
Number of Periods = 12*12 = 144
Rate per month = 8%/12

Monthly Payments or PMT = PV/(1-(1+r)-n)/r = 18000/(1-(1+8%/12)-144)/8%/12 = 194.84

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well